OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (anonymous):

i need help

OpenStudy (4everaddicted2anime):

What is the question?

OpenStudy (anonymous):

The graph shows the prices of different numbers of bushels of corn at a store in the current year. The table shows the prices of different numbers of bushels of corn at the same store for the previous year. A graph shows Number of Bushels on x-axis and Price of Corn in dollars on y-axis. The x-axis scale is shown from 0 to 14 at increments of 2, and the y-axis scale is shown from 0 to 98 at increments of 14. A straight line joins the ordered pairs 2, 14 and 4, 28 and 6, 42 and 8, 56 and 10, 70 and 12, 84. Previous Year Number of Bushels Price of Corn (dollars) 2 10 4 20 6 30 8 40 Part A: Describe in words how you can find the rate of change of a bushel of corn in the current year, and find the value. (5 points) Part B: How many dollars more is the price of a bushel of corn in the current year than the price of a bushel of corn in the previous year? Show your work. (5 points)

OpenStudy (anonymous):

it's extra credit work because i failed it :C

OpenStudy (anonymous):

can u help a llama out?

OpenStudy (anonymous):

@Dani_Rose

OpenStudy (4everaddicted2anime):

For part A all you need to do is find the slope

OpenStudy (anonymous):

owe okay

OpenStudy (anonymous):

can u help guide me i never been to great with graphs

OpenStudy (4everaddicted2anime):

To find the slope you don't need a graph. You are given coordinates that you can use.

OpenStudy (anonymous):

i know what the equations suppose to look like y = mx + b

OpenStudy (anonymous):

im just not sure what to do

OpenStudy (anonymous):

sorry for my supidity

OpenStudy (4everaddicted2anime):

The formula for slope is \[\frac{ y_2-y_1 }{ x_2-x_1 }=slope\]

OpenStudy (anonymous):

owe never knew that

OpenStudy (4everaddicted2anime):

Pick two of the point given to you and plug it into the equation

OpenStudy (anonymous):

ok 2,10 and 4,20

OpenStudy (4everaddicted2anime):

Okay now we have (2,10) and (4,20) Plug that into the equation to get\[\frac{ 20-10 }{ 4-2 }\] Simplify to get\[\frac{ 10 }{ 2 }\] 2 goes into 10 exactly 5 times so the slope would be 5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about part B?

OpenStudy (anonymous):

@4everaddicted2anime